1.     A car starts from rest and attains a speed of 40 m/s in 20 s. its average acceleration in m/s2 is



1
Expert's answer
2021-11-17T11:18:47-0500

"a=\frac{v-u}{t} \\

a=\frac{40m/s -0}{20s}\\

a=2\;m/s^2"